A, B, and C can finish a job in 6 days. If B and C work together, the job will take 9 days; if A and C work together, the job will take 8 days. In how many days can each man working alone do the job?
Answer by richwmiller(17219)

6/a+6/b+6/c=1
9/b+9/c=1
8/a+8/c=1
a = 18 days, b = 24 days , c = 72/5 = 14.4 days
check 1
6/18+6/24+30/72=1
24/72+18/72+30/72=1
ok 1
check 2
9/24+45/72=1
27/72+45/72=1
ok2
check3
8/18+40/72=1
32/72+40/72=1
ok3